The Mail1011 server is undergoing hardware maintenance.
There has been a failure detected with one of the CPUs serving this hardware node. Our VPS Administrators are bringing the node off line to swap out the affected CPU. UPDATE: Server replacement underway.
The iss1009 server has been slow today. Our Windows Administration team is working to diagnose the server and isolate the issue. We know the server is slow, but it is responsive at this time. FTP access is still available also. We will update the blog with more information as it comes in.
We were experiencing some issues with the web209 server. During a routine restart, it was found that the server motherboard is failing. The motherboard is being replaced, and we should have the server running soon. Expected ETA: 1 hour
Our systems administration department has detected a denial of service attack against the Shared IP address on web1016. In order to prevent the attack from compromising the server, they have blocked the Shared IP address for 8 hours.
We will be conducting system maintenance for our Cloud Environment on the morning of July 3 at 12:01AM EST (00:01 UTC-5). While the 6-hour maintenance window is scheduled through 06:00AM EST, we do not anticipate service outage for the entire period. During this maintenance period, a full shut down will occur. We will be conducting […]
VPS containers on the Windows based hardware node may be unavailable while the system admins attempt to correct possible file system errors.
The web server stopped responding to requests. At this time our System Administrators are attempting to stabilize and bring online the original web403. If those efforts fail, we will look to bringing the replacement server online as soon as it’s rsync is complete. Updates will follow as we get more information.
Web403 required a file system check after a routine restart and is temporarily down. After further investigation, the system admins discovered there may be a hardware issue with the server and are currently investigating the problem.
Our Systems Administration staff has detected an issue with the RAID connection on Mail402. At this point they have made some repairs to the mail server and after rebooting the server is undergoing a File System Check. Our current ETA is 60 -180 Minutes.
Our System Administrators have detected a Denial of Service attack against the shared IP address on webserver 915. In response, they have blocked the shared IP address for 4 hours to prevent the attack from impacting the rest of the server.
The LVZ1.opentransfer.com VPS Hardware Node was shut down for urgent maintenance. All services are up and running. Each VPS container is currently being rebooted one by one.
The mail server has become unresponsive. Our System Administrators are checking this mail server now. We will provide more information as it is available.
During a routine reboot the Web39.ixwebhosting.com server experienced some issues. We have swapped the server chassis on the machine, and it is currently running file system check to verify data integrity.
One of the hard drives in the RAID for MySQL504 has degraded. Backups have been created and the server is going down to replace the faulty drive. 6/14/2011 6:00PM EST – The server Raid rebuild has failed. Our system administration team is currently configuring a new server to replace the old one. The databases […]